OpenStudy (anonymous):

What is current in a gas discharge tube if 3.0 x 1017 electrons and 2.0 x 1016positive ions pass through a cross-section of the tube each second?

OpenStudy (muhammad_nauman_umair):

As, I=Q/t where, Q=ne here, n=3*10^17 and e=1.67*10^-11 and here t=1 As we know that current is the flow of electrons then we will consider only -ve particles not +ve particles then I=(3*10^17)*(1.67*10^-11)/1 answer is 5010000coulombs

OpenStudy (anonymous):

Thanks :)